Please I need today answer for This question and it is very important and I need solution for this issue with all the details , and help me with all the details.Please write your answer to me by typing, not by handwriting, so that I can read and understand your answer clearly.BR/Hassan.

1-Home has 600 unit of labor available.It can produce two goods, apples and nananas. the unit labour requirement in apple production is 4,while in banana production it is 2. Home is as described in problem 1.There is now also another country,foreign,with a labor force of 400. Foreign´s unit labor requirement in apple productionis 5, while in banana production it is 1.

a.-Graph foregin´s production possibility frontier.

b.- constuct the relative supply curve in the world.
Answer #1

4/2 = Labor requirement of apples in Home/Labor reequirement of bananas in Home

5/1 = Labor requirement of apple in Foreign/Labor requirement of bananas in Foreign

where Home lanbor = 600

Labor requirement for apple production = 4

Foreign Labor = 400

Labor requirement for bananas = 1

Hence (600/4)/400 = 0.375